OpenStudy (anonymous):

FInd the area of region enclosed by the given curves y=x+5 y^2=x y=2 y=-1 what I did was sliced it to parts. from o to -1 and 0 to 2 did I do it right? top half delta y and bottom delta x

OpenStudy (freckles):

hello are you here? I would start off with a drawing first

OpenStudy (freckles):

\[\int\limits_{-1}^{2} (\text{ \right }-\text{ \left } )dy\]

OpenStudy (anonymous):

yes I understand that and I wanted to split it in half at y=0 and take areas separately would that work? @freckles

OpenStudy (freckles):

yes but why do that?

OpenStudy (freckles):

seems much easier just to do one integral \[\int\limits_{-1}^2 (y^2-(y-5)) dy\]

OpenStudy (anonymous):

I guess I am just over thinking it. because today my teacher was talking about splitting integrals too for some equations

OpenStudy (freckles):

well you have a x=y^2 is right of x=y-5 so x=y^2 is greater than x=y-5

OpenStudy (freckles):